One mole of O2 gas is taken along an isochoric process from 100 Pa to 200 Pa at volume 1 m3 . Assuming the gas is ideal, find the change in the translational kinetic energy of a molecule. 


1
Expert's answer
2021-06-29T10:00:03-0400
"\Delta E=1.5kT\\\Delta E=1.5\frac{pV}{\nu N_a}
\\\Delta E=1.5\frac{(200-100)(1)}{1(6.02\cdot10^{23})}=2.5\cdot10^{-22}J"